OpenStudy (coledrf):

i give medals if you can help me please show work is x=-3 a solution to the equation -3x^2=-27

OpenStudy (anonymous):

well put -3 in for x and see if its true -3(-3)^2=-27 -3(9)=-27 -27=-27 thats true so yes it is a solution

OpenStudy (coledrf):

ok thank you